OpenStudy (anonymous):

When you divide both sides of an inequality by a positive value, you must change the direction of the inequality sign. True False

OpenStudy (anonymous):

Let's say we have 3x + 3 = 3 + 6. What does x equal to? In order for you to understand how signs behave/change, consider this method. Imagine that the simple numbers have to be on the right side of the equal and the ones with a variable (x in this case) will be on the left. Think that they go through the equal sign and automatically change signs. For example: 3x + 3 = 9; 3x = 9 - 3; 3x = 6; x = 6/2; x = 3 It's a lot more beautiful to do it this way until you get a good understanding...and then try to divide, etc. Simple rule of multiplication for signs: positive times negative equals to negative negative times negative equals to positive positive times positive equals positive This should help you figure out the answer. ;-)